# Synchronization of lasers with complex spatio-temporal dynamics

### v1

@@ B Comparison of master and slave laser
Among a batch of five nominally identical BA-VCSELs from the same manufacturer, a detailed characterization was performed on each device to identify the most suitable pair for the injection experiment. The selection was guided by two key criteria: similarity in birefringence frequency $\Delta\nu_{b}$ and in wavelength of the lasing modes. As the laser dynamics are strongly influenced by birefringence, we selected two devices exhibiting closely matched values, $\Delta\nu_{b}=8.8$ GHz for the master laser and $\Delta\nu_{b}=8.4$ GHz for the slave. Spectral alignment was also critical, since frequency detuning strongly affects the injection dynamics. Both lasers exhibited comparable emission wavelengths over the considered current range (222210101010 mA), ranging from $847.6$ to $848.6$ nm at a common temperature of 20°C.
@@
These analyses show that although the two selected VCSELs share key parameters such as birefringence and emission wavelength, their intrinsic dynamics are far from being stationary and furthermore differ significantly. This makes polarization-resolved injection experiments between them a challenging task. Once the pair of VCSELs was identified, polarization alignment in the injection path was ensured using a half-wave plate (see Fig. A4). Due to intrinsic birefringence and structural anisotropies, the $u$- and $v$-polarization axes of each VCSEL differ: for the master, $u=160^{\circ}$ and $v=70^{\circ}$; for the slave, $u=150^{\circ}$ and $v=60^{\circ}$ (note that these angles are given with respect to the horizontal axis of the laboratory frame, in contrast to the convention established in section A). Combined with the 45° rotation introduced by the optical isolator and further rotations from subsequent reflections, proper alignment of the injected field was achieved by tuning the HWP to ensure that the $u$-polarization of the master was injected into the $u$-axis of the slave to achieve polarization matching.
@@ C Dependency on laser parameters
Appendix 4 Injection experiments
@@ A Injection ratio
$\kappa_{\mathrm{inj}}=\frac{P_{\mathrm{inj}}}{P_{\mathrm{S,fr}}}\,.$ (A3)
The measured injection power accounts for optical losses due to the isolator, beam splitter, and other optical elements. Furthermore, the injection strength can be reduced via a variable neutral density filter (see Fig. A4), but it was set to minimal attenuation in the experiments presented here. $P_{\mathrm{S,fr}}$ is the free-running total output power of the slave laser, measured directly at the output, and includes both polarization components.
Figure A10: (a) Total output power $P_{\mathrm{S,fr}}$ of the slave laser in free-running conditions as a function of the pump current. The green (blue) dashed lines represent the injected power from the master laser for $I_{M}=8.8\,\mathrm{mA}$ ($I_{M}=6.17$ mA), respectively. (b) Corresponding injection strength $\kappa_{\mathrm{inj}}$ between the master and slave lasers for $I_{M}=6.17$ mA, plotted as a function of the slave laser’s pump current $I_{S}$.
In the first case of injection studied in the paper, the master laser operates at fixed parameters ($I_{M}=6.17\,\mathrm{mA}$) and delivers an input power of approximately $P_{\mathrm{inj}}=0.18$ mW at the entrance of the slave laser, indicated by the blue dashed line in Fig. A10(a). The output power of the slave $P_{\mathrm{S,fr}}$ varies nearly linearly with pump current as shown in Fig. A10, and this variation determines the resulting injection strength $\kappa_{\mathrm{inj}}$, plotted in Fig. A10(b). It is important to note that these values does not take into account reflections at the slave input facet or other factors affecting the coupling efficiency; they only reflect the available optical power at that point.
@@
In the second case of injection we study, at $I_{M}=8.8$ mA and $I_{S}=4.8$ mA, the injected power reaching the slave is approximately $0.37$ mW [green dashed line in Fig. A10(a)], while the total output power of the slave laser is $0.44$ mW, corresponding to an injection strength of $\kappa_{\mathrm{inj}}=0.84$.
However, it is important to note that the values of $\kappa_{\mathrm{inj}}$ presented here should be interpreted with caution. The measured master laser power corresponds to the value at the front of the slave laser. It is not possible to directly determine the portion that actually couples into the slave cavity. A rough estimate can nevertheless be made. In fact, assuming a typical front-facet reflectivity of $>90\%$, only a fraction of less than $10\%$ of the master power effectively enters the cavity. Similarly, the total output power of the slave laser used here does not reflect the real intracavity power, and is therefore underestimated. So while the injection strength is rather weak in our experiment, it allows us to rule out trivial synchronization effects driven purely by high-power injection.
@@ B Inverse synchronization
Figure A11: Spatio-spectral image of the master laser in $u-$polarization for $I_{M}$= 6.17 mA (bottom), with the spatio-spectral image of the slave laser at $\Delta\nu=0$ in $u-$ polarization (a) and at $\Delta\nu=1.7$ GHz in $v-$polarization (b).
The inverse synchronization phenomenon observed in the first injection case has been attributed to polarization-related dynamics. Figure A11 considers the first synchronization region and shows in panel (a) the case of $\Delta\nu=0$, where the $S_{u}(3,1)$ mode is spectrally aligned with the $M_{u}(3,1)$ mode. As a result, its intensity increases significantly, while neighboring modes are suppressed. This spectral alignment illustrates the strong mode interaction when their frequencies coincide. Figure A11(b) shows a similar effect, but in the orthogonal polarization of the slave laser ($v$-axis) at $\Delta\nu=1.7$ GHz: the $S_{v}(3,1)$ mode aligns with the $M_{u}(3,1)$ mode, leading to an intensity increase of $S_{v}(3,1)$ at the expense of adjacent modes. This alignment coincides with the occurrence of the negative correlation peak, suggesting anti-correlation between the $u$-polarized emission of the master and slave laser appears because the $v$-polarized emission of the slave is correlated with the master signal. This example illustrates that the excitation of a slave laser mode in the $v$-polarization systematically coincides with the occurrence of the inverse synchronization peaks.
@@ C Injection with variation of slave current and temperature
Figure A12: Correlation map between the master laser operated at $T_{M}=19^{\circ}$ and $I_{M}=9$ mA in the $u$-polarization, and the slave laser as a function of its temperature (horizontal axis) and injection current (vertical axis). The color scale represents the absolute value of correlation ($|C|\times 100$) of the unfiltered time traces. The black dashed lines follow the two principal series of correlation peaks.
Here, we present the evolution of the correlation between the master and slave lasers as a function of several experimentally scanned parameters. The master laser is held fixed at $I_{M}=9$ mA and $T_{M}=19^{\circ}$, while the slave laser is scanned in both temperature ($0.5^{\circ}$ step) and injection current ($0.01$ mA step). The time traces are measured as function of the pump current, following the same procedure used in the first injection case discussed earlier, while keeping the slave temperature fixed. Then the slave temperature is increased followed by another sweep of the current, covering a range of 16C to 20C.
Figure A12 displays the absolute value of the correlation of the unfiltered time traces as function of slave temperature and current. The temperature and current shifts are converted to the corresponding detuning, and the total detuning $\Delta\nu$ is the sum of the current and temperature detunings. The map reveals several series of correlation peaks, with the first and strongest one emerging at $\Delta\nu=0$ GHz. These correlation peak series appear as diagonals, reflecting conditions of spectral alignment between specific transverse modes of the master and slave lasers. The key implication of these diagonal series is that spectral alignment is a necessary condition for correlation, and can be achieved with both current and temperature variation. However, we also observe that the correlation varies along these diagonals, suggesting that other experimental parameters also influence the quality of synchronization.
@@ D Enhancement of Slave laser mode
Figure A13: Spatio-spectral images of the injected slave laser in $u$-polarization for different values of detuning, corresponding to the first injection case with $I_{M}=6.17\,\mathrm{mA}$. The case $\Delta\nu=0\,\mathrm{GHz}$ corresponds to the spectral alignment of the slave mode $S_{u}(3,1)$ with the master mode $M_{u}(3,1)$, while the case $\Delta\nu=80\,\mathrm{GHz}$ corresponds to the alignment of $S_{u}(2,1)$ with $M_{u}(0,1)$ (highlighted in green).
@@
The synchronization regions observed in the first injection case ($I_{M}=6.17\,\mathrm{mA}$) are closely associated with spectral alignment between a dominant transverse mode of the master laser and a transverse mode of the slave laser, depending on the frequency detuning $\Delta\nu$. Several distinct correlation peaks were identified, notably at $\Delta\nu=0\,\mathrm{GHz}$ and $\Delta\nu=80\,\mathrm{GHz}$, each corresponding to specific mode alignments. As shown in Fig. A13, at $\Delta\nu=0\,\mathrm{GHz}$, the slave mode $S_{u}(3,1)$ becomes spectrally aligned with the master mode $M_{u}(3,1)$. Under injection, this mode becomes more distinctly identifiable in the spatio-spectral image, indicating that it is effectively excited by the injected signal. This excitation is accompanied by a clear correlation peak, confirming the onset of synchronization. In contrast, for slightly detuned values such as $\Delta\nu=-1.1\,\mathrm{GHz}$ or $\Delta\nu=1.9\,\mathrm{GHz}$, the alignment is lost and the mode $S_{u}(3,1)$ becomes less distinguishable, which corresponds to a decrease in correlation.
@@
A similar behavior is observed near $\Delta\nu\approx 80\,\mathrm{GHz}$, where spectral alignment occurs between the slave mode $S_{u}(2,1)$ and the master mode $M_{u}(0,1)$. The slave mode again becomes more clearly visible and identified, indicating its excitation through injection, and this results in a secondary synchronization peak. These observations confirm that the emergence of a well-defined transverse mode in the slave laser, driven by spectral alignment with a dominant mode of the master, plays a central role in enabling synchronization, even without matching the spatial profiles.

### v2

@@ B Comparison of master and slave laser
Among a batch of five nominally identical BA-VCSELs from the same manufacturer, a detailed characterization was performed on each device to identify the most suitable pair for the injection experiment. The selection was guided by two key criteria: similarity in birefringence frequency $\Delta\nu_{b}$ and in wavelength of the lasing modes. As the laser dynamics are strongly influenced by birefringence, we selected two devices exhibiting closely matched values, $\Delta\nu_{b}=8.8$ GHz for the master laser and $\Delta\nu_{b}=8.4$ GHz for the slave. Spectral alignment was also critical, since frequency detuning strongly affects the injection dynamics. Both lasers exhibited comparable emission wavelengths over the considered current range (221010 mA), ranging from $847.6$ to $848.6$ nm at a common temperature of 20°C.
@@
These analyses show that although the two selected VCSELs share key parameters such as birefringence and emission wavelength, their intrinsic dynamics are far from being stationary and furthermore differ significantly. This makes polarization-resolved injection experiments between them a challenging task. Once the pair of VCSELs was identified, polarization alignment in the injection path was ensured using a half-wave plate (see Fig. S1). Due to intrinsic birefringence and structural anisotropies, the $u$- and $v$-polarization axes of each VCSEL differ: for the master, $u=160^{\circ}$ and $v=70^{\circ}$; for the slave, $u=150^{\circ}$ and $v=60^{\circ}$ (note that these angles are given with respect to the horizontal axis of the laboratory frame, in contrast to the convention established in section A). Combined with the 45° rotation introduced by the optical isolator and further rotations from subsequent reflections, proper alignment of the injected field was achieved by tuning the HWP to ensure that the $u$-polarization of the master was injected into the $u$-axis of the slave to achieve polarization matching.
Furthermore, the estimated relaxation oscillation frequencies range from 0.7 to 6.7 GHz for the master laser and from 0.6 to 7.9 GHz for the slave laser as a function of pump current, but no significant influence of their matching on the synchronization quality is observed in our experiments.
@@ C Dependency on laser parameters
Supplementary Section 3 Injection experiments
@@ A Injection ratio
$\kappa_{\mathrm{inj}}=\frac{P_{\mathrm{inj}}}{P_{\mathrm{S,fr}}}\,$ (S1)
The measured injection power accounts for optical losses due to the isolator, beam splitter, and other optical elements. Furthermore, the injection strength can be reduced via a variable neutral density filter (see Fig. S1), but it was set to minimal attenuation in the experiments presented here. $P_{\mathrm{S,fr}}$ is the free-running total output power of the slave laser, measured directly at the output, and includes both polarization components.
Figure S7: (a) Total output power $P_{\mathrm{S,fr}}$ of the slave laser in free-running conditions as a function of the pump current. The green (blue) dashed lines represent the injected power from the master laser for $I_{M}=8.8\,\mathrm{mA}$ ($I_{M}=6.17$ mA), respectively. (b) Corresponding injection strength $\kappa_{\mathrm{inj}}$ between the master and slave lasers for $I_{M}=6.17$ mA, plotted as a function of the slave laser’s pump current $I_{S}$.
In the first case of injection studied in the paper, the master laser operates at fixed parameters ($I_{M}=6.17\,\mathrm{mA}$) and delivers an input power of approximately $P_{\mathrm{inj}}=0.18$ mW at the entrance of the slave laser, indicated by the blue dashed line in Fig. S7(a). The output power of the slave $P_{\mathrm{S,fr}}$ varies nearly linearly with pump current as shown in Fig. S7, and this variation determines the resulting injection strength $\kappa_{\mathrm{inj}}$, plotted in Fig. S7(b). It is important to note that these values do not take into account reflections at the slave input facet or other factors affecting the coupling efficiency; they only reflect the available optical power at that point.
@@
In the second case of injection we study, at $I_{M}=8.8$ mA and $I_{S}=4.8$ mA, the injected power reaching the slave is approximately $0.37$ mW [green dashed line in Fig. S7(a)], while the total output power of the slave laser is $0.44$ mW, corresponding to an injection strength of $\kappa_{\mathrm{inj}}=0.84$.
However, it is important to note that the values of $\kappa_{\mathrm{inj}}$ presented here should be interpreted with caution. The measured master laser power corresponds to the value at the front of the slave laser. It is not possible to directly determine the portion that actually couples into the slave cavity. A rough estimate can nevertheless be made. In fact, assuming a typical front-facet reflectivity of $90\%$, only a fraction of less than $10\%$ of the master power effectively enters the cavity. Similarly, the total output power of the slave laser used here does not reflect the real intracavity power, and is therefore underestimated. So while the injection strength is rather weak in our experiment, it allows us to rule out trivial synchronization effects driven purely by high-power injection.
@@ B Inverse synchronization
Figure S8: Spatio-spectral image of the master laser in $u$-polarization for $I_{M}$= 6.17 mA (bottom), with the spatio-spectral image of the slave laser at $\Delta\nu=0$ in $u$- polarization (a) and at $\Delta\nu=1.7$ GHz in $v$-polarization (b).
The inverse synchronization phenomenon observed in the first injection case has been attributed to polarization-related dynamics. Figure S8 considers the first synchronization region and shows in panel (a) the case of $\Delta\nu=0$, where the $S_{u}(3,1)$ mode is spectrally aligned with the $M_{u}(3,1)$ mode. As a result, its intensity increases significantly, while neighboring modes are suppressed. This spectral alignment illustrates the strong mode interaction when their frequencies coincide. Figure S8(b) shows a similar effect, but in the orthogonal polarization of the slave laser ($v$-axis) at $\Delta\nu=1.7$ GHz: the $S_{v}(3,1)$ mode aligns with the $M_{u}(3,1)$ mode, leading to an intensity increase of $S_{v}(3,1)$ at the expense of adjacent modes. This alignment coincides with the occurrence of the negative correlation peak, suggesting anti-correlation between the $u$-polarized emission of the master and slave laser appears because the $v$-polarized emission of the slave is correlated with the master signal. This example illustrates that the excitation of a slave laser mode in the $v$-polarization systematically coincides with the occurrence of the inverse synchronization peaks.
@@ C Injection with variation of slave current and temperature
Figure S9: Correlation map between the master laser operated at $T_{M}=19^{\circ}$C and $I_{M}=9$ mA in the $u$-polarization, and the slave laser as a function of its temperature (horizontal axis) and injection current (vertical axis). The color scale represents the absolute value of correlation ($|C|\times 100$) of the unfiltered time traces. The black dashed lines follow the two principal series of correlation peaks.
Here, we present the evolution of the correlation between the master and slave lasers as a function of several experimentally scanned parameters. The master laser is held fixed at $I_{M}=9$ mA and $T_{M}=19^{\circ}$C, while the slave laser is scanned in both temperature ($0.5^{\circ}$ step) and injection current ($0.01$ mA step). The time traces are measured as function of the pump current, following the same procedure used in the first injection case discussed earlier, while keeping the slave temperature fixed. Then the slave temperature is increased followed by another sweep of the current, covering a range of 16C to 20C.
Figure S9 displays the absolute value of the correlation of the unfiltered time traces as function of slave temperature and current. The temperature and current shifts are converted to the corresponding detuning, and the total detuning $\Delta\nu$ is the sum of the current and temperature detunings. The map reveals several series of correlation peaks, with the first and strongest one emerging at $\Delta\nu=0$ GHz. These correlation peak series appear as diagonals, reflecting conditions of spectral alignment between specific transverse modes of the master and slave lasers. The key implication of these diagonal series is that spectral alignment is a necessary condition for correlation, and can be achieved with both current and temperature variation. However, we also observe that the correlation varies along these diagonals, suggesting that other experimental parameters also influence the quality of synchronization.
@@ D Enhancement of Slave laser mode
Figure S10: Spatio-spectral images of the injected slave laser in $u$-polarization for different values of the detuning $\Delta\nu$ for the first injection case with $I_{M}=6.17\,\mathrm{mA}$. (a) At $\Delta\nu=0\,\mathrm{GHz}$ we find the spectral alignment of the slave mode $S_{u}(3,1)$ with the master mode $M_{u}(3,1)$, (b) and at $\Delta\nu=80\,\mathrm{GHz}$ the alignment of $S_{u}(2,1)$ with $M_{u}(0,1)$. The spectrally aligned slave modes are highlighted by the blue rectangles.
@@
The synchronization regions observed in the first injection case ($I_{M}=6.17\,\mathrm{mA}$) are closely associated with spectral alignment between a strong transverse mode of the master laser and a transverse mode of the slave laser as function of the frequency detuning $\Delta\nu$, see Fig. 2. Several distinct correlation peaks were identified, notably at $\Delta\nu=0\,\mathrm{GHz}$ and $\Delta\nu=80\,\mathrm{GHz}$, each corresponding to the alignment of specific lasing modes. When the slave mode $S_{u}(3,1)$ becomes spectrally aligned with the master mode $M_{u}(3,1)$ at $\Delta\nu=0\,\mathrm{GHz}$, the intensity of $S_{u}(3,1)$ increases significantly while that of neighboring modes decreases due to the injection as shown in Fig. S10(a). This selective excitation of $S_{u}(3,1)$ is accompanied by a clear correlation peak, confirming the onset of synchronization. In contrast, for slightly detuned values such as $\Delta\nu=-1.1\,\mathrm{GHz}$ or $\Delta\nu=1.9\,\mathrm{GHz}$, the spectral alignment is lost and the mode $S_{u}(3,1)$ is less intense, which goes along with a decrease in correlation.
@@
A similar behavior is observed near $\Delta\nu\approx 80\,\mathrm{GHz}$, where spectral alignment occurs between the slave mode $S_{u}(2,1)$ and the master mode $M_{u}(0,1)$. The intensity of the slave mode increases like in the first example, indicating its enhanced excitation due to spectrally aligned injection, and this results in a secondary synchronization peak. These observations demonstrate that the enhancement of a certain transverse mode of the slave laser, driven by spectral alignment with a strong mode of the master, plays a central role in enabling synchronization, but matching the spatial profiles of the involved master and slave modes is not required.
E Correlation evolution with low- and high-pass filters
Figure S11: (a,b,c1) RF-spectra of the master (black) and slave (red) lasers. Evolution of the correlation as a function of cutoff frequency for low-pass (a,b,c2) and high-pass (a,b,c3) filtering for three detuning values: (a) $\Delta\nu=0\,\mathrm{GHz}$, (b) $\Delta\nu=44\,\mathrm{GHz}$, and (c) $\Delta\nu=115\,\mathrm{GHz}$. (d) Correlation as a function of detuning using a low-pass filter at 1 GHz cutoff and (e) a high-pass filter at 1 GHz.
The preferential synchronization of low-frequency components appears to depend on the system and on the specific synchronization region. As shown in Fig. 2(d) of the main article, the correlation value increases as the high-frequency components are progressively removed, thereby revealing one of the characteristic timescales that favors higher correlation value.
Here, we present the evolution of the correlation values between the master and slave lasers for three correlation peaks, associated with the second injection experiment shown in Fig. 3. We determine how the correlation evolves when applying a low- or high-pass filter, with the corresponding results displayed in Fig. S11.
These measurements of the evolution of the correlation value under low-pass and high-pass filtering show that, for the first two correlation peaks (a) and (b), the correlation increases as the high-frequency components are removed (Fig. S11 ($a_{2}$), ($b_{2}$)). The region corresponding to the third peak (c), at a detuning of 115 GHz, exhibits a lower synchronization level, associated with faster timescales. This ability to easily synchronize low-frequency components is consistent with the literature [S8-S11] and with the majority of our experiments.